How much power is required to charge a capacitor of 9.4uF to 5.4x10^2 V in 48 seconds?
Answer #1

here,

capacitance , C = 9.4 uF = 9.4 * 10^-6 F

potential difference , V = 540 V

time period , T = 48 s

the power required , P = energy stored /T

P = (0.5 * C * V^2)/T

P = (0.5 * 9.4 * 10^-6 * 540^2) /48 W

P = 2.86 * 10^-2 W

the power required is 2.86 * 10^-2 W